• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:02:25 on August 25, 2001

    Entry number 72445

    keyword=Beam Time Accounting
    Hall A

    8/25/2001 0800-1600
    Experiment #: e97103
    Shift Leader: Simon Sirca
    Shift Workers: Piotr Zolnierczuk, Francesco Cusanno

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    23 50 54 25 53 35 40 51 5.51666666666667
    Beam Available
    but Not in Use
    37 5 0 30 0 0 0 0 1.2
    Beam Not Available
    or Unacceptable
    0 5 6 5 7 25 20 9 1.28333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    23 50 1 30 1 1 1 1 6.71666666666667
    Planned Config-
    uration Change
    37 0 0 30 0 0 0 0 1.11666666666667
    Unplanned
    Experiment Down
    0 10 0 0 0 0 0 00.166666666666667
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Ac 0 10 0 0 0 0 0 0 0.166666666666667
    Other 0 0 0 0 0 0 0 0 0

    Comments: Continued kinematics #3 (beam OUT, target IN, holding field 90 deg) until about 11:30, then switching to (target OUT, holding field 270 deg), ready at about 12:00. Halog not functional for about 1.5 hours, minor DAQ problems. Beam off for about 3/4 hour due to RF zone down.

    Data Entry by: Simon Sirca